If w is the complex cube root of unity then the value of (1+w^2)^3

Grade:12th pass

1 Answers

Priyansh Gupta
2 years ago
we know 1+w +w^2 =0
Therefore 1+w^2 =-w 
 
Substituting in the equation we get (-w)^3 which is equal to – (w)^3 = -1
since w^3 =1